OpenStudy (anonymous):

Which numbers should be multiplied to obtain 183^2 -139^2? A. 44 and 44 B. 1936 and 1936 C. 44 and 322 D. 33350 and 33628

hartnn (hartnn):

use the famous formula \(\Large a^2-b^2 = (a+b)(a-b)\) here, a = 183 b = 139

hartnn (hartnn):

what is a+b =...? a-b =... ?